A boat travels 60km upstream in 5 hours. The boat travels the same distance downstream in 3 hours. What is the rate of the boat in still water? What is the rate of the current?
Answer by mananth(16949)

Boat speed =x mph
current speed =y mph
against current x-y 6.00 hours
With current x+y 3.00 hours

Distance = same= 60 miles
t=d/r
60 / ( x - y )= 6.00
6 ( x - y ) = 60.00
6 x -6 y = 60 ....................1
60 / ( x + y )= 3.00
3.00 ( x + y ) = 60
3.00 x + 3.00 y = 60 ...............2
Multiply (1) by 1.00
Multiply (2) by 2.00
we get
6 x + -6 y = 60
6 x + 6 y = 120
12 x = 180
/ 12
x = 15 mph

plug value of x in (1)
6 x -6 y = 60
90 -6 y = 60
-6 y = 60 -90
-6 y = -30
y = 5 mph
Boat 15 mph
current 5 mph
m.ananth@hotmail.ca